# Diffusion-Aided Bandwidth-Efficient Semantic Communication with Adaptive Requests

### v2

@@ I Introduction
Traditional communication systems aim to reliably transmit bits under channel constraints. In many scenarios, however, what ultimately matters is the utility of the received content but not exact bitwise accuracy [sagduyu2023task]. This has motivated semantic communication, in which a transmitter conveys a representation tailored to preserve task-relevant meaning, with system performance evaluated in the task space [liu2024survey, wang2025diffusion]. Such systems can be implemented with end-to-end neural network approaches that train encoders and decoders simultaneously, with task model optionally integrated [liu2024explainable, grassucci2023generative]. For image transmission, sending pixels demands heavy transmission bandwidth while compact semantics sustain coherent reconstructions even at low bitrate [tung2025multi].
@@
Generative models enhance this paradigm by providing a strong prior at the receiver. Latent-based diffusion models, for instance, can synthesize high-quality images using caption and/or latent as inputs, which allows transmitting latents rather than pixels to reduce bandwidth [blattmann2023align, rombach2022high]. However, simply replacing pixels with latents does not fully resolve the rate bottleneck since latents still consume a significant portion of the bit budget, and text-only conditioning often results in reconstructions that diverge from the source content [cicchetti2024language, wang2025trustworthy]. This motivates us to propose a two-stage approach to further improve bandwidth efficiency, that is, sending a random subset of latent blocks and the caption in the first pass, and then employing feedback to request additional blocks only when necessary to improve semantic fidelity. A specially designed latent diffusion model can effectively leverage such sparse evidence by inpainting missing regions and produce a globally coherent image that aligns with available clues and the intended semantics [xie2023smartbrush], while language guidance further steers generation toward the desired content with minimal overhead [manukyan2023hd].
@@
This strategy naturally requires a feedback mechanism that decides when to request more blocks and when to stop. Classical hybrid ARQ uses decoder checks or CRC to trigger incremental redundancy and pursues bit-level reliability, but in semantic communication the receiver does not have the ground-truth image and bitwise correctness is not the objective, so a conventional ACK is neither available nor appropriate. Recent studies on semantic HARQ either retain a bit-level ACK and use semantic features mainly for combining or proxy triggers [zhou2022adaptive, liu2024adaptive], or move the decision to the task space using feature-based or topology-based detectors [fei2025topology, li2025semantic, sheng2025semantic]. In this paper, we address this gap with a receiver-driven semantic acknowledgment. After each reconstruction the receiver performs caption-guided latent diffusion inpainting and generates a caption, and then compares it with the intended description using ROUGE-L (a language-based score). If the score meets a threshold the receiver issues a semantic ACK and stops transmission, otherwise it requests a small set of additional latent blocks from a reserved pool and fuses them through inpainting. This places both the trigger and the stopping rule in the semantic space and adapts the transmitted payload via feedback. It differs from prior work by using a language metric as the control signal and by coupling selective latent requests with diffusion-based inpainting, which yields substantial bandwidth savings over pixel-centric communication and full-latent decoding while maintaining stable semantic fidelity and strong visual quality.
@@ II-A Transmitter-side Design
The roles of $z$ and $y$ are asymmetric since the caption is short but semantically decisive, whereas the latent carries most of the visual detail. In the experiments we simply apply the same lightweight channel code to both streams, so that we can vary the SNR to emulate different link qualities while keeping the receiver-side reconstruction logic unchanged. Token-level noise can sharply distort meaning [zhang2020spelling], while latent tensors are continuous and decoded by a smooth generator, so small residual distortions are visually milder than discrete-token errors [rombach2022high]. Noise-aware latent handling with channel denoising diffusion models [wu2024cddm] is left to future work.

### v3

@@ I Introduction
Traditional communication systems aim to reliably transmit bits under channel constraints. In many scenarios, however, what ultimately matters is the utility of the received content but not exact bitwise accuracy [12]. This has motivated semantic communication, in which a transmitter conveys a representation tailored to preserve task-relevant meaning, with system performance evaluated in the task space [9, 17]. Such systems can be implemented with end-to-end neural network approaches that train encoders and decoders simultaneously, with task model optionally integrated [7, 4]. For image transmission, sending pixels demands heavy transmission bandwidth while compact semantics sustain coherent reconstructions even at low bitrate [15].
@@
Generative models enhance this paradigm by providing a strong prior at the receiver. Latent-based diffusion models, for instance, can synthesize high-quality images using caption and/or latent as inputs, which allows transmitting latents rather than pixels to reduce bandwidth [1, 11]. However, simply replacing pixels with latents does not fully resolve the rate bottleneck since latents still consume a significant portion of the bit budget, and text-only conditioning often results in reconstructions that diverge from the source content [2, 16]. This motivates us to propose a two-stage approach to further improve bandwidth efficiency, that is, sending a random subset of latent blocks and the caption in the first pass, and then employing feedback to request additional blocks only when necessary to improve semantic fidelity. A specially designed latent diffusion model can effectively leverage such sparse evidence by inpainting missing regions and produce a globally coherent image that aligns with available clues and the intended semantics [19], while language guidance further steers generation toward the desired content with minimal overhead [10].
@@
This strategy naturally requires a feedback mechanism that decides when to request more blocks and when to stop. Classical hybrid ARQ uses decoder checks or CRC to trigger incremental redundancy and pursues bit-level reliability, but in semantic communication the receiver does not have the ground-truth image and bitwise correctness is not the objective, so a conventional ACK is neither available nor appropriate. Recent studies on semantic HARQ either retain a bit-level ACK and use semantic features mainly for combining or proxy triggers [21, 8], or move the decision to the task space using feature-based or topology-based detectors [3, 5, 13]. In this paper, we address this gap with a receiver-driven semantic acknowledgment. After each reconstruction the receiver performs caption-guided latent diffusion inpainting and generates a caption, and then compares it with the intended description using ROUGE-L (a language-based score). If the score meets a threshold the receiver issues a semantic ACK and stops transmission, otherwise it requests a small set of additional latent blocks from a reserved pool and fuses them through inpainting. This places both the trigger and the stopping rule in the semantic space and adapts the transmitted payload via feedback. It differs from prior work by using a language metric as the control signal and by coupling selective latent requests with diffusion-based inpainting, which yields substantial bandwidth savings over pixel-centric communication and full-latent decoding while maintaining stable semantic fidelity and strong visual quality.
@@ II-A Transmitter-side Design
The roles of $z$ and $y$ are asymmetric since the caption is short but semantically decisive, whereas the latent carries most of the visual detail. In the experiments we simply apply the same lightweight channel code to both streams, so that we can vary the SNR to emulate different link qualities while keeping the receiver-side reconstruction logic unchanged. Token-level noise can sharply distort meaning [20], while latent tensors are continuous and decoded by a smooth generator, so small residual distortions are visually milder than discrete-token errors [11]. Noise-aware latent handling with channel denoising diffusion models [18] is left to future work.